OpenStudy (anonymous):

the equilibrium mixture given beow contain 5 moles of CO and 2 moles of H2O the volume of container is 1 litre Co+H2O-------->Co2+H2 at equilibrium 0.5 moles of H2O are Present the value of Kc for the reaction is=?

OpenStudy (mathmate):

First check/balance equation: CO + H2O = CO2 + H2 Volume of container = 1 L so all quantities in moles are equivalent to concentration in mol/L Use ICE table CO H2O CO2 H2 Initial 5 2 0 0 change -x -x x x Final 5-x 2-x x x All reactants and products are in gas form: At equilibrium, 0.5 moles of H2O left, so 0.5=2-x, or x=1.5 \(Kc=\large \frac{[CO2]^1[H2]^1}{[CO]^1[H2p]^1}\) \(\large =\frac{x^2}{(5-x)(2-x)}\) Substitute x=1.5 to find Kc

OpenStudy (anonymous):

thanx

OpenStudy (mathmate):

You're welcome! :)